There are two treasure chests filled with gemstones. Treasure chest A contains: 6
rubies, 9 emeralds, and 5 diamonds. Treasure chest B contains: 3 rubies, 7 emeralds,
and 10 diamonds. A diamond is selected from one of the chests at random. What is the
probability that the diamond was selected from Treasure chest A?
Answer #1

Solution:
Contingency table can be written as follows

Treasure Chests

Rubies

Emeralds

Diamonds

A

6

9

5

20

B

2

7

10

19

8

16

15

39

We need to find
P(Chest A | Diamond) here we will use conditional probability
P(Chest A| Diamond) = P(ChestA and Diamond)/ P(Diamond) = (5/39) / (15/39) = 5/15 = 0.33